Jo is thinking of a number. 4 less than 9 times the number is 176. Find Jo's number.
Answer by ewatrrr(24785)


Hi,
Let x represent the number
Write as we Read***
9x - 4 = 176
9x = 180
x = 20
checking our answer
9*20-4 = 180- 4 = 176